# BayesianFitForecast: A User-Friendly R Toolbox for Parameter Estimation and Forecasting with Ordinary Differential Equations

### v2

@@ 3   Results and Discussion
    calibrationperiods <- c(17)    forecastinghorizon <- 10    vars <- c("S", "E", "I", "R", "C")    ode_system <- ’      diff_var1 = -params1 * vars3 * vars1 / params5      diff_var2 = params1 * vars3 * vars1 / params5 - params3 * vars2      diff_var3 = params3 * vars2 - params2 * vars3      diff_var4 = params2 * vars3      diff_var5 = params4 * params3 * vars2’    paramsfix <- c(0,1,1,1,1)    fitting_index <- 5    fitting_diff <- 1    errstrc <- 3    params1_prior <- "uniform(0, 10)"    params2_prior <- 1/4.1    params3_prior <- 1/1.9    params4_prior <- 1    params5_prior <- 550,000    niter <- 1000We retain all other parameters from the existing file. Special attention is given to the paramsfix option, where $\rho$ is fixed, and the prior is set accordingly. Since params4_prior = 1 implies $\rho=1$, there is no need to exclude this parameter from the model.
Next, open run_MCMC to execute the MCMC algorithm. Begin by modifying line 5 in run_MCMC to to reflect the new options file: options_SEIR_sanfrancisco_Ex1.R. Select all lines of code, then click ‘Run’ in RStudio. This process may take some time to complete. Once finished, an .Rdata file will be generated in the same directory, indicating that the data is ready for analysis. Proceed by opening the run_analyzeResults.R script to generate the results. Ensure that you modify the name of the options file here as well. Once this section is executed, the results will be produced. Relevant figures and tables are shown in Figure 5 and Table 2 and 3.
Figure 5: Bayesian fitting of the SEIR model to the 1918 influenza dataset from San Francisco was performed using a Poisson error structure, calibrated over 17 days. The model captures the dynamics of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are held constant, while the prior distribution for $\beta$ is uniform(0,10). All parameters were constrained to have non-negative values, with a total population size fixed at $550,000$. Initial conditions were set as (549996, 0, 4, 0, 4), based on the first recorded data point. The MCMC algorithm was executed with 1,000 iterations across two independent chains to ensure robust convergence. In the top panel, the red curve represents the model’s best fit to the data, while the black dashed lines delineate the $95\%$ prediction interval (PI). The circles represent the observed data points for comparison.The bottom panel presents a histogram of the posterior distributions for the transmission rate parameter $\beta$, and the reproduction number $R\_0$, derived from the frequency of MCMC iterations. The median, mean, and $95\%$ credible intervals (CI) are highlighted, giving a comprehensive summary of the parameter estimates and their uncertainties.
@@
  • The model fits the data to the number of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are held constant, with a uniform prior distribution $(0,10)$ for $\beta$. All parameters are constrained to have non-negative values. The population size is fixed at 550,000, with initial conditions set to (549996, 0, 4, 0, 4), based on the first recorded data point. The MCMC algorithm was executed with 1,000 iterations across two chains to ensure robust convergence.
@@
  • The model is fitted to the number of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are treated as constants, while the prior distribution for $\beta$ is uniform(0,10). All parameters are bounded below by zero to ensure non-negative values. The population size is set at $550,000$, with initial conditions of (549996, 0, 4, 0, 4) based on the first recorded data point in the dataset. The MCMC algorithm was executed for 1,000 iterations across two chains, ensuring robust posterior estimates.
@@
Figure 6: The Bayesian fitting of the SEIR model to the first 17 days of the 1918 influenza pandemic in San Francisco was performed using a negative binomial error structure. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5), with all parameters having a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains. The top panel shows the model fit to the data, with the red curve representing the median of fitted curves and black dashed lines indicating the 95% PI, while the circles depict the actual data points. The bottom panels show the histograms of $\beta$ (left), $\phi$ (center), and $R_{0}$ (right).
@@
  • The model is fitted over 17 days to newly infected individuals, represented by $\frac{dC}{dt}$. Parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants. The prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters are constrained to non-negative values. The population size is set to 550,000, with initial conditions (549996, 0, 4, 0, 4) based on the first data point. The MCMC algorithm used 1,000 iterations across two chains.
@@
  • The model is fitted over 17 days to newly infected individuals, represented by $\frac{dC}{dt}$. Parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants. The prior distribution for $\beta$ is uniform(0,10), and for $\phi$, it is exponential(5). All parameters are bounded below by zero. The population size is set to 550,000, with initial conditions of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm used 1,000 iterations across two chains.
@@
Bayesian, SF 1918 flu data, Poisson error structure 15.24 45.35 88.24 3.33
Bayesian, SF 1918 flu data, negative binomial error structure 5.76 68.24 100 3.57
  • The model is fitted over 17 days to newly infected individuals, represented by $\frac{dC}{dt}$. Parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with initial conditions (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm used 1,000 iterations across two chains. The red curve represents the median of fitted curves, while the black dashed lines indicate the 95% confidence interval, and circles depict actual data points.
@@ 3.3 Generating Forecasts with the Best Model
In this example, we present the best model for forecasting 10 days after calibrating 17 days of the SF 1918 flu dataset. During the calibration period, the Poisson error structure performs better in terms of MAE, MSE, and WIS metrics, while the negative binomial model excels in $95\%$ PI coverage. However, the negative binomial model demonstrates superior performance across all metrics in the forecasting period. Figure 7 provide a forecasting for both Poisson and negative binomial error structure. Also, a comparison of the performance metrics is given in Table 7. The related option files for this practice example are ”options_SEIR_sanfrancisco_Ex3_Negbin” and ”options_SEIR_sanfrancisco_Ex3_Poisson”.
Figure 7: Forecasting results for the SEIR model, calibrated over 17 days with a 10-day forecast horizon. This figure allows for a direct comparison, highlighting differences in forecast accuracy and uncertainty between the two error structures. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains. The red curve represents the median of a series of fitted curves, while the black dashed lines indicate the 95% PI. The circles depict the actual data points. The left panel uses a negative binomial error structure while the right one has a Poisson error structure.
@@
  • The table highlights differences in MAE, MSE, WIS, and 95% PI coverage, showing the superiority of the negative binomial model in forecasting performance and uncertainty quantification over a 10-day horizon, using a 17-day calibration period. The models are fitted to newly infected individuals, represented by $\frac{dC}{dt}$. Parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants, with prior distributions for $\beta$ as uniform(0,10) and $\phi$ as exponential(5). All parameters are bounded below by zero. The population size is 550,000, with initial conditions (549996, 0, 4, 0, 4), based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains.
@@ 3.4 Comparison of the SEIR Model with the Exponential Growth Model
Table 8: A comparison of performance metrics between the SEIR model and the exponential growth model using Bayesian fitting to the San Francisco 1918 flu dataset with negative binomial error structure, using 17 days of calibration data with a 10-day forecasting horizon.
@@
  • The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. In the SEIR model, parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants. Prior distributions for $r$ and $\beta$ are uniform(0,10), and for $\phi$ is exponential(5). All parameters are bounded below by zero. The population size is 550,000, with initial conditions (549996, 0, 4, 0, 4) for the SEIR model and (4) for the exponential growth model, based on the first recorded data point. The MCMC algorithm ran with 1,000 iterations across two chains.
@@ 3.5 The SEIR Model with Estimating the Initial number of infected people
Figure 9: Bayesian fitting of the SEIR model to the initial 17 days of the 1918 influenza pandemic in San Francisco using a normal error structure. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{1.4}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\sigma$ is cauchy(0, 2.5), with all parameters having a lower bound of zero. The population size is 550,000, with an initial condition of (N-$i_{0}$, 0, $i_{0}$, 0, $i_{0}$) with the purpose of estimating the first recorded data point $i_{0}$. The prior distribution for $i_{0}$ is normal(0,10) by assuming an upper bound of 100. The MCMC algorithm was run with 1,000 iterations across two chains. The top panel shows the model fit to the data, with the red curve representing the median of fitted curves and black dashed lines indicating the 95% PI, while the circles depict the actual data points. The bottom panels show the histograms of $\beta$ (left), $\sigma$ (center), and $i_{0}$ (right).
@@ 3.6 Fitting the SEIR model to multiple data sets using the simulated data
In this final practice example, we simulate data using the forward solution of the SEIR model (17) with the parameters $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and $N=100{,}000$ and adding normally distributed noise with $SD=5$. The Excel file containing the simulated data can be found in the toolbox with the name "curve-SEIR_plain_I-dRdt-dCdt-dist1-0-factor1-5". In this file, the first column represents the days, the second column represents the number of infectious individuals ($I$), the third column represents the derivative of recovered individuals ($\frac{dR}{dt}$), and the fourth column represents the derivative of cumulative cases ($\frac{dC}{dt}$). We consider three different modeling scenarios: (i) fitting the SEIR model to all the data, (ii) fitting the model to only the infectious individuals ($I$) and the derivative of cumulative cases ($\frac{dC}{dt}$), and (iii) fitting the model to only the derivative of cumulative cases ($\frac{dC}{dt}$). A calibration period of 70 days and a forecasting horizon of 30 days were used. Figure 10 shows the fitted curves of the newly incidence together with the prediction bands from the three scenarios. As expected, the results from the first scenario, where all available data is used, outperform those from the second scenario, which only uses the $I$ and $\frac{dC}{dt}$ columns. The second scenario, in turn, performs better than the third scenario, which relies solely on $\frac{dC}{dt}$. This is because including more data provides more information for model fitting. Table 9 highlights an interesting observation: even though the third scenario performs better in the calibration phase, it underperforms in the forecasting phase. The parameter estimates from the third scenario are greatly different from the true values, and the first scenario provides more accurate parameter estimates (Table 10). This implies that there might be an overfitting to the calibration period in the third scenario which has less data, and this leads to poor forecasting. Using multiple states provides more information and better avoids the overfitting. The option files used for each of these scenarios are:"options_SEIR_simulated_Ex6.R", "options_SEIR_simulated_Ex6_I_dcdt.R", and "options_SEIR_simulated_Ex6_
@@
Figure 10: We fit the SEIR model to the simulated data generated by the forward solution of the SEIR model  (17) using the parameters $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and $N=100{,}000$, with added noise sampled from $N(0,25)$, using three different scenarios: (i) left panel—fitting the SEIR model to all the data, (ii) middle panel—fitting the SEIR model to only the first $I$ and the third column $\frac{dC}{dt}$, and (iii) right panel—fitting the SEIR model to only the third column $\frac{dC}{dt}$. In all scenarios, the figures display the time series of newly infected individuals, represented by $\frac{dC}{dt}$.
@@
(i)$I$, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 30 3.06 15.59 96.67 1.99
(ii)$I$, $\frac{dC}{dt}$ 30 3.22 16.02 93.33 2.02
(iii)$\frac{dC}{dt}$ 30 3.04 14.08 96.67 1.90
@@
(i)$I$, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 70 100.98 17216.70 87.14 65.78
(ii)$I$, $\frac{dC}{dt}$ 70 131.93 35634.15 84.29 93.13
(iii)$\frac{dC}{dt}$ 70 431.25 464089.54 60.0 287.49
@@
  • The model utilized the following parameters: transmission rate $\beta=0.5$, recovery rate $\gamma=0.25$, progression rate $\kappa=1$, reporting rate $\rho=1$, and a total population size of $N=100{,}000$. Three distinct scenarios were analyzed: (i) fitting the SEIR model to the entire dataset, (ii) fitting the model to both the infected population $I$ and the rate of new cases, $\frac{dC}{dt}$, and (iii) fitting the SEIR model solely to the rate of new cases, $\frac{dC}{dt}$.
@@
Table 10: Parameter estimates obtained by fitting the SEIR model to 30 days of simulated data.
Scenario Forecasting Horizon $\beta$ $\gamma$ $\kappa$
(i) $I$, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 70 0.47 (0.4, 0.92) 0.24 (0.21, 0.27) 1.09 (0.16, 9.07)
(ii) $I$, $\frac{dC}{dt}$ 30 0.57 (0.38, 5.28) 0.21 (0.16, 0.25) 0.37 (0.02, 7.81)
(iii) $\frac{dC}{dt}$ 30 6.29 (0.44, 9.83) 3.98 (0.25, 7.78) 0.45 (0.09, 5.85)
@@
  • Parameter estimation was performed by fitting the SEIR model to simulated data generated using the forward solution of the SEIR model (17) with parameters $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and a population size of $N=100{,}000$. Three scenarios were considered: (i) fitting the model to all compartments, (ii) fitting to $I$ and $\frac{dC}{dt}$, and (iii) fitting solely to $\frac{dC}{dt}$.

### v3

@@ 3 Results and Discussion
    calibrationperiods <- c(17)    forecastinghorizon <- 10    vars <- c("S", "E", "I", "R", "C")    ode_system <- ’      diff_var1 = -params1 * vars3 * vars1 / params5      diff_var2 = params1 * vars3 * vars1 / params5 - params3 * vars2      diff_var3 = params3 * vars2 - params2 * vars3      diff_var4 = params2 * vars3      diff_var5 = params4 * params3 * vars2’    paramsfix <- c(0,1,1,1,1)    fitting_index <- 5    fitting_diff <- 1    errstrc <- 3    params1_prior <- "uniform(0, 10)"    params2_prior <- 1/4.1    params3_prior <- 1/1.9    params4_prior <- 1    params5_prior <- 550000    niter <- 1000We retain all other parameters from the existing file. Special attention is given to the paramsfix option, where $\rho$ is fixed, and the prior is set accordingly. Since params4_prior = 1 implies $\rho=1$, there is no need to exclude this parameter from the model.
Next, open run_MCMC to execute the MCMC algorithm. Begin by modifying line 5 in run_MCMC to reflect the new options file: options_SEIR_sanfrancisco_Ex1.R. Select all lines of code, then click ‘Run’ in RStudio. This process may take some time to complete. Once finished, an .Rdata file will be generated in the same directory, indicating that the data is ready for analysis. Proceed by opening the run_analyzeResults.R script to generate the results. Ensure that you modify the name of the options file here as well. Once this section is executed, the results will be produced. Relevant figures and tables are shown in Figure 6 and Table 2 and 3.
Figure 6: Bayesian fitting of the SEIR model to the 1918 influenza dataset from San Francisco was performed using a Poisson error structure, calibrated over 17 days. The model captures the dynamics of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are held constant, while the prior distribution for $\beta$ is uniform(0,10). All parameters were constrained to have non-negative values, with a total population size fixed at $550,000$. Initial conditions were set as (549996, 0, 4, 0, 4), based on the first recorded data point. The MCMC algorithm was executed with 1,000 iterations across two independent chains to ensure robust convergence. In the top panel, the red curve represents the model’s best fit to the data, while the black dashed lines delineate the $95\%$ prediction interval (PI). The circles represent the observed data points for comparison.The bottom panel presents a histogram of the posterior distributions for the transmission rate parameter $\beta$, and the reproduction number R_0R\_0, derived from the frequency of MCMC iterations. The median, mean, and $95\%$ credible intervals (CI) are highlighted, giving a comprehensive summary of the parameter estimates and their uncertainties.
@@
00footnotetext: The model fits the data to the number of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are held constant, with a uniform prior distribution $(0,10)$ for $\beta$. All parameters are constrained to have non-negative values. The population size is fixed at 550,000, with initial conditions set to (549996, 0, 4, 0, 4), based on the first recorded data point. The MCMC algorithm was executed with 1,000 iterations across two chains to ensure robust convergence.
@@
00footnotetext: The model is fitted to the number of newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are treated as constants, while the prior distribution for $\beta$ is uniform(0,10). All parameters are bounded below by zero to ensure non-negative values. The population size is set at $550,000$, with initial conditions of (549996, 0, 4, 0, 4) based on the first recorded data point in the dataset. The MCMC algorithm was executed for 1,000 iterations across two chains, ensuring robust posterior estimates.
@@
Figure 7: The Bayesian fitting of the SEIR model to the first 17 days of the 1918 influenza pandemic in San Francisco was performed using a negative binomial error structure. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5), with all parameters having a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains. The top panel shows the model fit to the data, with the red curve representing the median of fitted curves and black dashed lines indicating the 95% PI, while the circles depict the actual data points. The bottom panels show the histograms of $\beta$ (left), $\phi$ (center), and $R_{0}$ (right).
@@
00footnotetext: The model is fitted over 17-days to newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains.
@@
00footnotetext: The model is fitted over 17-days to newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains.
@@
Bayesian, SF 1918 flu data set, Poisson error structure 15.24 45.35 88.24 3.33
Bayesian, SF 1918 flu data set, negative binomial error structure 5.76 68.24 100 3.57
00footnotetext: The model is fitted over 17-days to newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains. The red curve represents the median of a series of fitted curves, while the black dashed lines indicate the 95% confidence interval. The circles depict the actual data points.
@@ 3.3 Generating Forecasts with the Best Model
In this example, we present the best model for forecasting 10 days after calibrating 17 days of the SF 1918 flu dataset. During the calibration period, the Poisson error structure performs better in terms of MAE, MSE, and WIS metrics, while the negative binomial model excels in $95\%$ PI coverage. However, the negative binomial model demonstrates superior performance across all metrics in the forecasting period. Figure 8 provide a forecasting for both Poisson and negative binomial error structure. Also, a comparison of the performance metrics is given in Table 7. The related option files for this practice example are "options_SEIR_sanfrancisco_Ex3_Negbin" and "options_SEIR_sanfrancisco_Ex3_Poisson".
Figure 8: Forecasting results for the SEIR model, calibrated over 17 days with a 10-day forecast horizon. This figure allows for a direct comparison, highlighting differences in forecast accuracy and uncertainty between the two error structures. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains. The red curve represents the median of a series of fitted curves, while the black dashed lines indicate the 95% PI. The circles depict the actual data points. The left panel uses a negative binomial error structure while the right one has a Poisson error structure.
@@
00footnotetext: The table highlights differences in MAE, MSE, WIS, and $95\%$ PI coverage, emphasizing the superiority of the negative binomial model in forecasting performance and uncertainty quantification over a 10-day horizon, using a 17-day calibration period. The models are fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains.
@@ 3.4 Comparison of the SEIR Model with the Exponential Growth Model
Table 8: A comparison of performance metrics between the SEIR model and the exponential growth model using Bayesian fitting to the San Francisco 1918 flu dataset with negative binomial error structure, using 17 days of calibration data with a forecasting horizon 10.
@@
00footnotetext: The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. In the SEIR model, the parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants. The prior distribution for $r$ and $\beta$ is uniform(0,10), and for $\phi$ is exponential(5). All parameters have a lower bound of zero. The population size is 550,000, with an initial condition of (549996, 0, 4, 0, 4) for the SEIR model, and (4) for the exponential growth model based on the first recorded data point. The MCMC algorithm was run with 1,000 iterations across two chains.
@@ 3.5 The SEIR Model with Estimating the Initial number of infected people
Figure 10: Bayesian fitting of the SEIR model to the initial 17 days of the 1918 influenza pandemic in San Francisco using a normal error structure. The model is fitted to the newly infected individuals, represented by $\frac{dC}{dt}$. The parameters $\kappa=\frac{1}{1.9}$, $\gamma=\frac{1}{4.1}$, and $\rho=1$ are constants, while the prior distribution for $\beta$ is uniform(0,10) and for $\sigma$ is cauchy(0, 2.5), with all parameters having a lower bound of zero. The population size is 550,000, with an initial condition of (N-$i_{0}$, 0, $i_{0}$, 0, $i_{0}$) with the purpose of estimating the first recorded data point $i_{0}$. The prior distribution for $i_{0}$ is normal(0,10) by assuming an upper bound of 100. The MCMC algorithm was run with 1,000 iterations across two chains. The top panel shows the model fit to the data, with the red curve representing the median of fitted curves and black dashed lines indicating the 95% PI, while the circles depict the actual data points. The bottom panels show the histograms of $\beta$ (left), $\sigma$ (center), and $i_{0}$ (right).
@@ 3.6 Fitting the SEIR model to multiple data sets using the simulated data
In this final practice example, we simulate data using the forward solution of the SEIR model (16) with the parameters $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and N=100, [^0] N=100{,}000 and adding normally distributed noise with $SD=5$. The Excel file containing the simulated data can be found in the toolbox with the name "curve-SEIR_plain_I-dRdt-dCdt-dist1-0-factor1-5". In this file, the first column represents the days, the second column represents the number of infectious individuals ($I$), the third column represents the derivative of recovered individuals ($\frac{dR}{dt}$), and the fourth column represents the derivative of cumulative cases ($\frac{dC}{dt}$). We consider three different modeling scenarios: (i) fitting the SEIR model to all the data, (ii) fitting the model to only the infectious individuals ($I$) and the derivative of cumulative cases ($\frac{dC}{dt}$), and (iii) fitting the model to only the derivative of cumulative cases ($\frac{dC}{dt}$). A calibration period of 70 days and a forecasting horizon of 30 days were used. Figure 11 shows the fitted curves of the newly incidence together with the prediction bands from the three scenarios. As expected, the results from the first scenario, where all available data is used, outperform those from the second scenario, which only uses the $I$ and $\frac{dC}{dt}$ columns. The second scenario, in turn, performs better than the third scenario, which relies solely on $\frac{dC}{dt}$. This is because including more data provides more information for model fitting. Table 9 highlights an interesting observation: even though the third scenario performs better in the calibration phase, it underperforms in the forecasting phase. The parameter estimates from the third scenario are greatly different from the true values, and the first scenario provides more accurate parameter estimates (Table 10). This implies that there might be an overfitting to the calibration period in the third scenario which has less data, and this leads to poor forecasting. Using multiple states provides more information and better avoids the overfitting. The option files used for each of these scenarios are:"options_SEIR_simulated_Ex6.R", "options_SEIR_simulated_Ex6_I_dcdt.R", and "options_SEIR_simulated_Ex6_
@@
Figure 11: We fit the SEIR model to the simulated data generated by the forward solution of the SEIR model  (16) using the parameters $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and N=100, [^0] N=100{,}000, with added noise sampled from $N(0,25)$, using three different scenarios: (i) left panel—fitting the SEIR model to all the data, (ii) middle panel—fitting the SEIR model to only the first II and the third column $\frac{dC}{dt}$, and (iii) right panel—fitting the SEIR model to only the third column $\frac{dC}{dt}$. In all scenarios, the figures display the time series of newly infected individuals, represented by $\frac{dC}{dt}$.
@@
(i)II, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 30 3.08 15.65 96.67 1.99
(ii)II, $\frac{dC}{dt}$ 30 3.23 16.11 93.33 2.03
[…]
@@
(i)II, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 70 98.86 16487.18 87.14 65.43
(ii)II, $\frac{dC}{dt}$ 70 131.26 35876.48 84.29 92.26
[…]
@@
00footnotetext: The model utilized the following parameters: transmission rate $\beta=0.5$, recovery rate $\gamma=0.25$, progression rate $\kappa=1$, reporting rate $\rho=1$, and a total population size of N=100, [^0] N=100{,}000. Three distinct scenarios were analyzed: (i) fitting the SEIR model to the entire dataset, (ii) fitting the model to both the infected population II and the third column of data, $\frac{dC}{dt}$, and (iii) fitting the SEIR model solely to the third column, $\frac{dC}{dt}$.
@@
Table 10: Parameter estimates obtained by fitting the SEIR model to 30-days of simulated data.
Scenario Calibration $\beta$ $\gamma$ $\kappa$
(i) $I$, $\frac{dR}{dt}$, $\frac{dC}{dt}$ 30 0.48 (0.4, 0.83) 0.24 (0.21,0.27) 1.08 (0.19,8.84)
(ii)II, $\frac{dC}{dt}$ 30 0.56 (0.38,4.88) 0.21 (0.16,0.25) 0.38 (0.02,7.82)
[…]
@@
00footnotetext: Parameter estimation was performed by fitting the SEIR model to simulated data generated using the forward solution of the SEIR model  (16) with the following parameters: $\beta=0.5$, $\gamma=0.25$, $\kappa=1$, $\rho=1$, and a population size of N=100, [^0] N=100{,}000. Three different scenarios were considered: (i) fitting the SEIR model to the entire dataset, (ii) fitting the SEIR model to only the first column (II) and the third column $\frac{dC}{dt}$, and (iii) fitting the model solely to the third column $\frac{dC}{dt}$.
3.7 Fitting the SEIR model with a time-dependent parameter to the simulated data
In this practice example, we simulate data using the forward solution of the SEIR model (16) with the time-dependent parameter $\beta$, defined as below:
$\beta=\beta(t)=\begin{cases}\beta_{0},&\text{if }t<t_{\text{int}}\\ \beta_{1}+(\beta_{0}-\beta_{1})e^{-q(t-t_{\text{int}})},&\text{if }t\geq t_{\text{int}}\end{cases}$ (19)
This piecewise function describes a transition in $\beta(t)$ from an initial value $\beta_{0}$ to a new value $\beta_{1}$ after a specific intervention time $t_{\text{int}}$. For times $t$ earlier than $t_{\text{int}}$, the function remains constant at $\beta_{0}$. After $t_{\text{int}}$, the function decays exponentially towards $\beta_{1}$, with the rate of decay governed by the parameter $q$.
We generate the forward solution of this model over the time period $[0,100]$, using the following parameters: $\beta_{0}=0.5$, $\beta_{1}=0.1$, $q=0.1$, $t_{\text{int}}=10$, $\gamma=\frac{1}{7}$, $\kappa=\frac{1}{5}$, $\rho=0.6$, and N=1,000, [^0] N=1{,}000{,}000. The Excel file containing the simulated data is located in the toolbox under the name "seir_simulated". In this file, the first column represents the days, and the second column contains the derivative of the cumulative cases, denoted by $\frac{dC}{dt}$. The option file related to this practice example is under the name "options_seir_timedep_Ex7.R" located in the toolbox directory.
Figure 12 illustrates an excellent fit of the SEIR model with a time-dependent transmission rate to the synthetic data. Notably, the impact of the time-dependent transmission rate is evident, as the epidemic curve is asymmetric with respect to the peak, reflecting the dynamic nature of the transmission over time.
Table 11 shows that the estimated parameters do not exactly match the true values, but they are reasonably close. This suggests that while parameter estimates may not be fully reliable when using a time-dependent parameter in the model, the model can still provide a good fit to the data.
Figure 12: Fitting of the SEIR model (16) with a time-dependent transmission rate (19) to synthetic data generated from the forward simulation of the same model, using parameters $\beta_{0}=0.5$, $\beta_{1}=0.1$, $q=0.1$, $t_{\text{int}}=10$, $\gamma=\frac{1}{7}$, $\kappa=\frac{1}{5}$, $\rho=0.6$, and N=1,000, [^0] N=1{,}000{,}000.
Table 11: Convergence and posterior statistics for the SEIR model with a time-dependent transmission rate fitted to synthetic data.
Parameter Mean Median CI95 Neff
$\beta_{0}$ 0.91 0.67 (0.55, 1.62) 26.08 1.01
$\beta_{1}$ 0.34 0.06 (0.02, 1.20) 25.77 1.01
$q$ 0.16 0.10 (0.10, 0.45) 33.43 1.00
$\kappa$ 0.11 0.12 (0.08, 0.14) 30.95 1.01
$\gamma$ 0.14 0.11 (0.07, 0.25) 28.97 1.00
$\rho$ 0.45 0.48 (0.33, 0.55) 29.37 1.00
00footnotetext: The model was fitted to synthetic epidemic data generated from the SEIR model with a time-dependent transmission rate. Parameters $\beta_{0}$, $\beta_{1}$, $q$, $\kappa$, $\gamma$, and $\rho$ were estimated using MCMC, with 10,000 iterations across two chains. The effective sample size ($N_{\text{eff}}$) and potential scale reduction factor ($\hat{R}$) are reported to assess convergence.